I need help on the quadratic equation: 2x(to the 2square)-7x+5=0? How do you solve it?
Answer by kiru_khandelwal(79)

2x^2 -7x +5 = 0
=> 2x^2 -2x - 5x + 5 =0
=> 2x(x-1) -5(x-1) = 0
=> (2x-5)(x-1) = 0
=> According to the zero product rule
either 2x-5 =0 => x = 5/2
or x-1 = 0 => x =1
So, x= 5/2 or 1